# Internet of Mirrors for Connected Healthcare and Beauty: A Prospective Vision

### v1

@@ Abstract
Internet of Things, Assistive Technologies, Smart Mirrors, Beauty Tech.
I Introduction
@@
The emergence of the IoT has significantly supported the shift towards increased interconnectivity and automation in many sectors such as healthcare, transportation, and smart cities [1], [2]. The increasing number of ‘things’, internet-connected embedded systems, has opened possibilities for numerous everyday physical objects to interact with each other and achieve shared goals such as enhanced energy efficiency, improved health monitoring, and streamlined supply chains [3, 4, 5] In recent years, in homes, offices and even industrial contexts, there has been an upsurge of smart environments, appliances and devices that all connect wirelessly over the internet or through local networks. The increased deployment of IoT has led to the development of various interconnected frameworks that fall under the umbrella of IoT such as the Internet of Vehicles (IoV) [6] and the Internet of Musical Things (IoMusT) [7]. There has been an increased research focus in multiple domains such as transportation and smart cities but there is minimal academic research on the application of IoT technologies in the cosmetics and beauty industries. Comparatively, there are ongoing significant research efforts for IoT in healthcare with the rise of telemedicine. However, there is still a need for enhanced and seamless connectivity to expand the reach and accessibility of healthcare systems to the public and enable advanced use-cases. This will not only reduce healthcare staff workload but also improve the patient experience on a holistic level to eventually support better healthcare outcomes [8]. Additionally, this research focus often relates to applications such as emergency management, post-discharge care through health monitoring and patient authentication [9] but there is a notable lack of academic attention given to specialised and elective healthcare services like dermatology, dentistry, and optometry. Many aspects of elective healthcare services overlap with the cosmetics industry indicating the potential for a unified platform that could encompass both sectors.
There is an increasing prevalence and demand for AI-driven services and personalisation in the cosmetics and beauty industry with the integration of technologies like IoT and Augmented Reality (AR) to provide services such as personalised product recommendations and cosmetic colour customisers [10]. This demand for personalisation is highlighted in an Accenture survey [11] which found that 75% of consumers are more likely to make a purchase when recommendations are tailored to them. Similarly, a Forrester study [12] found that 77% of consumers have chosen, recommended, or paid extra for a brand that offers a personalised service or experience. The current capabilities of commercially available smart beauty products, often termed beauty tech, are limited to information delivery and product customisations with services like portable skin analysers, cosmetic curation apps and skin analysis apps [13]. Although these products provide personalised user experiences, they are not widespread and accessible and serve only a single purpose of service. This indicates the need for a multi-purpose unified platform that can offer personalised services. This issue extends into elective healthcare or aesthetic medicine as well in areas such as dentistry and optometry. In these areas although there are smart services such as diagnostic support [14] and mobile app-integrated smart medical instruments [15], these also only serve a single purpose and do not provide the clinician or patient with a holistic view of the treatment process.
The nature of personalised services in the cosmetics and beauty industry and the elective healthcare sector are visual. An existing smart system that has the potential to support these services is a smart mirror. A smart mirror refers to an intelligent interface that operates as a mirror and can interact with users and display various responses and information. Currently, smart mirror products are becoming increasingly available for assistive beauty and assistive living applications. The common features of most smart mirrors include the display of multimedia data and the option for the user to interact with contextual information such as weather conditions and interaction with options such as voice recognition or a touch screen dashboard [16]. Within beauty tech, there is an increasing market for smart mirrors that offer results-driven services such as virtual makeup try-on or product recommendations [17]. However, drawbacks of beauty tech smart mirrors have been found in consumer interviews conducted by Yum et al [18] which revealed three main limitations, namely; lack of personalisation to individual/consumer needs; lack of research on data exchange between smart devices and the smart mirror (i.e. the IoT environment); and limited AR features such as real-time virtual try-on. In general, regardless of the smart mirror application, whether it is for the smart home environment or beauty tech, smart mirror products are currently limited to stand-alone systems.
To the best of our knowledge, there is no existing vision or concept of interconnected IoT-based smart mirrors. Capitalising on this, this article aims to provide a unified vision for an interconnected system of mirrors coined as the Internet of Mirrors (IoM) under the IoT framework. This system will be conceptualised with the existing concept of smart mirrors to support assistive technologies, primarily the health and beauty sector. We will focus on defining the IoM ecosystem as well as discussing the state-of-the-art technologies that will enable the existence of the IoM. This article seeks to bridge the gaps between an implementation-focused approach to smart mirrors and the need for interconnected services in the assistive health and beauty sectors. The gap between the industrial development of smart mirrors and beauty tech and the corresponding lack of development in this domain in academia will also be highlighted.
@@ II-A Smart Mirrors
Table I: Examples of 5G and LPWAN integration in smart mirrors.
Wireless Technology
Feature
Explanation
5G
Cloud Connectivity
Connection with fifth-generation (5G) optimised cloud services would allow smart mirrors to access and store data efficiently and quickly.
App Integration
Built-in 5G compatible apps can utilise 5G’s high speeds and low latency communication for the delivery of high quality services and analytics.
WiFi Hotspot
The smart mirrors can act as wireless access points which would allow other devices to connect to 5G networks through them.
LPWANs
Asset Tracking
Tracking of the location of physical assets and other personal items would allow users to locate their items quickly and easily.
Energy Management
Integration with LPWAN-enabled energy management systems would enable smart mirrors to monitor and adjust energy consumption in real time.
Smart Home Integration
Integration with smart home systems that use LPWANs would enable smart mirrors to communicate with other devices in the home, such as smart locks, smart thermostats, and security systems.
Remote Monitoring
Users could monitor their mirror’s usage and performance remotely to allow for ease in maintenance and troubleshooting.
@@
With the growing commercial smart mirror market [24], a system of interconnected mirrors could enable many new opportunities not only for the cosmetics and health industries but also for other domains such as smart homes and smart education. The visualisation aspect of the smart mirror interface also has the potential to transform user accessibility for consumers and patients by offering a more enhanced and innovative health and beauty experience.
II-B Connectivity
The connectivity of the IoM ecosystem is fundamental to its operation and development. To support the envisioned multi-directional communication between smart mirrors and from smart mirror to the user (and vice versa), appropriate network protocols and standards must be considered. Depending on the application and domain, different local and remote wireless communication protocols can be deployed. The connectivity of the IoM would facilitate the processing of computationally heavy commands and requests to off-load the computation from the embedded micro-controller in the smart mirror. Although the applications of the IoM in the cosmetics industry are not time critical in comparison to health applications, real-time use cases such as remote consultancy would still require the network to fulfil low latency, high reliability and synchronisation requirements. The connectivity element of the IoM will not only enable interconnection between IoM smart mirrors and devices but could also facilitate additional services to support the user experience. Table I provides an overview of two key wireless technologies, 5G and Low Power Wide Area Networks (LPWANs), and possible features that could be enabled by them in the IoM.
@@ II-B Connectivity
II-C Applications and Services
Whether the application domain for the IoM is assistive health, beauty or living, there are several services that can be developed from the sensed and communicated data from the smart mirrors and their connectivity features. These services can target consumers (beauty and cosmetic customers, patients), professionals (cosmeticians, dermatologists, doctors) and service providers (brands, stores, clinics). To establish an efficient interactive interface of the smart mirrors, these services will need to consider real-time computation and requirements. Fig.4 shows the possible remote and co-located interactions between different users in the IoM ecosystem.
III Enabling Technologies
In this section, a list of state-of-the-art technologies related to the IoM is presented. We aim to describe the results of various application domains that lead to the conceptualisation of the IoM. A top-level overview of the presented technologies and their potential integration in the IoM is shown at the end of this section in Table III.
III-A Standalone Smart Mirrors
@@
In recent years, there have been some increased research efforts within this domain to develop and integrate new features such as the use of Infrared (IR) sensors for user presence detection [34], gesture control and tracking for music control [35] as well as WiFi connectivity [36] for mobile synchronisation with the digitally displayed data [37]. There have also been efforts in the development of plugins and web APIs to support smart mirror applications [38] [39]. There has also been an increase in the number of commercially available smart mirrors for both assistive beauty and assistive living applications. There are many commercial options of stand-alone mirrors for assistive beauty purposes as well as assistive health. Commonly available smart mirrors with their main features have been shown in Table II.
Table II: Commercially available smart mirrors and their features.
@@
Sephora Mirror [40]
QR code generation for in-store and online makeup, skincare and fragrance recommendations
HiMirror [41]
Adjustable lighting, skin analysis, product recommendation, cosmetics tutorial
MemoMi [42]
Image capturing and storage,AR eyewear and clothes, Smartphone and SNS sharing
Charlotte Tilbury MagicMirror [43]
Augmented makeup, importing images
Shiseido Digital Counseling Mirror [44]
Product information, application instructions, skin check, smartphone compatibility
Panasonic New Magic Mirror [45]
Skin condition identification, virtual try-on
Modiface [46]
Product recommendation, virtual try-on
Lumini Kiosk V2 [47]
Skin analysis, product recommendation, skin comparison with other users
Icon AI Z Mirror [48]
Adjustable lighting, skin analysis, product recommendation, virtual try-on
Morror [49]
Adjustable lighting, cosmetics tutorial, weather information
Simplehuman Sensor Mirror Pro [50]
Smart lighting
Kohler Verdera Voice Lighted Mirror [51]
Voice activation integrated with Alexa and Google assistant
iHome Reflect Smart Vanity Mirror [52]
Lighting and speaker control with smartphone
Capstone Connected Home Smart [53]
Integration with other smart home devices, built-in speaker, voice activation
@@
Figure 5: A simple IoT-enabled smart mirror schematic.
@@
III-B Skin Imaging Methodologies
@@
III-C Real-Time Health Monitoring

### v2

# Internet-of-Mirrors (IoM) for Connected Healthcare and Beauty: A Prospective Vision

@@ Abstract
Internet of Things , Assistive Technologies , Smart Mirrors , Beauty Tech
1 Introduction
1.1 Background
The emergence of the Internet-of-Things (IoT) has significantly supported the shift towards increased interconnectivity and automation in many sectors such as healthcare, transportation, and smart cities [1], [2]. The increasing number of ‘things’, internet-connected embedded systems, has opened possibilities for numerous everyday physical objects to interact with each other and achieve shared goals such as enhanced energy efficiency, improved health monitoring, and streamlined supply chains [3, 4, 5].
In recent years, in homes, offices and even industrial contexts, there has been an upsurge of smart environments, appliances and devices that all connect wirelessly over the internet or through local networks. The increased deployment of IoT has led to the development of various interconnected frameworks that fall under the umbrella of IoT such as the Internet-of-Vehicles (IoV) [6] and the Internet-of-Musical-Things (IoMusT) [7]. There has been an increased research focus in multiple domains such as transportation and smart cities but there is minimal academic research on the application of IoT technologies in the cosmetics and beauty industries. Comparatively, there are ongoing significant research efforts for IoT in healthcare with the rise of telemedicine. However, there is still a need for enhanced and seamless connectivity to expand the reach and accessibility of healthcare systems to the public and enable advanced use-cases. This will not only reduce healthcare staff workload but also improve the patient experience on a holistic level to eventually support better healthcare outcomes [8]. Additionally, this research focus often relates to applications such as emergency management, post-discharge care through health monitoring and patient authentication [9] but there is a notable lack of academic attention given to specialised and elective healthcare services like dermatology, dentistry, and optometry. Many aspects of elective healthcare services overlap with the cosmetics industry indicating the potential for a unified platform that could encompass both sectors.
The nature of personalised services in the cosmetics and beauty industry and the elective healthcare sector are visual. An existing smart system that has the potential to support these services is a smart mirror. A smart mirror refers to an intelligent interface that operates as a mirror and can interact with users and display various responses and information. Currently, smart mirror products are becoming increasingly available for assistive beauty and assistive living applications as evidenced by their growing market presence and adoption across sectors such as retail and healthcare [10]. The common features of most smart mirrors include the display of multimedia data and the option for the user to interact with contextual information such as weather conditions and interaction with options such as voice recognition or a touch screen dashboard [11]. However, despite these features, current smart mirror products are limited to standalone systems, lacking the interconnectivity and comprehensive integration that characterise more mature IoT ecosystems. This gap in connectivity, particularly in the realms of health and beauty, presents an opportunity for innovative solutions that can leverage the visual nature of these services while harnessing the full potential of IoT technologies.
1.2 Motivation
There is an increasing prevalence and demand for AI-driven services and personalisation in the cosmetics and beauty industry with the integration of technologies like IoT and Augmented Reality (AR) to provide services such as personalised product recommendations and cosmetic colour customisers [12]. This demand for personalisation is highlighted in an Accenture survey [13] which found that 75% of consumers are more likely to make a purchase when recommendations are tailored to them. Similarly, a Forrester study [14] found that 77% of consumers have chosen, recommended, or paid extra for a brand that offers a personalised service or experience. This trend towards personalisation aligns with the broader healthcare industry’s recognition of the critical need for comprehensive, patient-tailored services. As highlighted in [15], personalised care that considers patients’ unique characteristics, including risk factors and clinical history, can significantly improve treatment efficiency and patient satisfaction.
Building upon this trend, the beauty and healthcare industries have been quick to adopt new technologies. However, the current landscape of smart beauty and health technologies faces several critical challenges. The current capabilities of commercially available smart beauty products, often termed beauty tech, are limited to information delivery and product customisations with services like portable skin analysers, cosmetic curation apps and skin analysis apps [16]. Although these products provide personalised user experiences, they are not widespread and accessible and serve only a single purpose of service. This indicates the need for a multi-purpose unified platform that can offer personalised services. This fragmentation extends into elective healthcare or aesthetic medicine as well in areas such as dentistry and optometry [17]. In these areas although there are smart services such as diagnostic support [18] and mobile app-integrated smart medical instruments [19], these also only serve a single purpose and do not provide the clinician or patient with a holistic view of the treatment process.
Within beauty tech, smart mirrors have emerged as a promising technology, offering results-driven services such as virtual makeup try-on or product recommendations [20]. However, drawbacks of beauty tech smart mirrors have been found in consumer interviews conducted in [21] three main limitations are highlighted, namely; lack of personalisation to individual/consumer needs; lack of research on data exchange between smart devices and the smart mirror (i.e. the IoT environment); and limited AR features such as real-time virtual try-on. In general, regardless of the smart mirror application, whether it is for the smart home environment or beauty tech, smart mirror products are currently limited to stand-alone systems, failing to integrate with broader IoT ecosystems.
These limitations mirror the challenges faced in the broader healthcare sector. As highlighted in [15], the fragmentation of medical data and the difficulty of generating comprehensive results across different populations hinder the delivery of truly personalised care. Furthermore, while IoT sensors and AI technologies offer the potential for continuous, real-time data collection and analysis, deploying these technologies at scale while ensuring data security and privacy remains a significant challenge. The convergence of challenges in both beauty and healthcare sectors highlights the need for a unified, multi-purpose platform offering advanced personalisation beyond current standalone systems. This platform should integrate data holistically from multiple sources, provide context-aware suggestions, and offer cross-domain insights considering the interplay between health and appearance. It should enable continuous, non-invasive monitoring and seamless data integration while ensuring robust privacy protections. By addressing these aspects, such a platform would offer a more nuanced, adaptive user experience that evolves with individual needs, overcoming the limitations of current fragmented systems in both health and beauty sectors.
1.3 Contribution
To address these challenges and capitalise on the potential of a unified platform for interconnected smart technologies in the health and beauty sectors, we propose a novel concept: the Internet-of-Mirrors (IoM). To the best of our knowledge, there is no existing vision or concept of interconnected IoT-based smart mirrors. Capitalising on this, this article aims to provide a unified vision for an interconnected system of mirrors coined as the IoM under the IoT framework. This system will be conceptualised with the existing concept of smart mirrors to support assistive technologies, primarily the health and beauty sector. We will focus on defining the IoM ecosystem as well as discussing the state-of-the-art technologies that will enable the existence of the IoM. This article seeks to bridge the gaps between an implementation-focused approach to smart mirrors and the need for interconnected services in the assistive health and beauty sectors. Although this paper builds on existing smart mirrors and IoT technologies, it provides a comprehensive and integrated perspective on the inter-operation of all ecosystem components, addressing the gaps in personalised services for the cosmetics and healthcare industries. This perspective is aimed at bridging the gap between the industrial development of smart mirrors and personalised services and the corresponding lack of development in academia, especially in the domains of elective healthcare and beauty tech.
Overall, the IoM is envisioned to address the unique needs and challenges of the cosmetics and elective healthcare sectors which require the seamless integration of diverse data types to provide personalised and interactive services, which is not typically targeted by IoT frameworks. The IoM differentiates itself by focusing on the interoperability and interconnectivity of smart mirrors within an ecosystem, allowing for enhanced data exchange, real-time analysis, and personalised recommendations. This level of integration is crucial for delivering comprehensive and effective assistive services, ensuring that users receive tailored and holistic care. The user experience directly dictates the methods and approaches that will lead to the realisation of the IoM and this requirement of user-centric criticality also differentiates the IoM from other IoT sub frameworks. Additionally, the IoM can support everyday living, self-care, and non-invasive healthcare by incorporating typically non-smart parts of user lives into the smart city vision, making it distinct from existing smart mirror technologies with limited capabilities.
@@ 2.1 Smart Mirrors
With the increasing interest in existing smart mirrors [27], the integration of the envisioned IoM could enable many new opportunities not only for the cosmetics and health industries but also for other domains such as smart homes and smart education. The visualisation aspect of the smart mirror interface also has the potential to transform user accessibility for consumers and patients by offering a more enhanced and innovative health and beauty experience.
2.2 Connectivity
Table 1: Examples of 5G and LPWAN integration in smart mirrors.
Wireless Technology Feature Explanation
5G Cloud Connectivity Connection with fifth-generation (5G) optimised cloud services would allow smart mirrors to access and store data efficiently and quickly.
App Integration Built-in 5G compatible apps can utilise 5G’s high speeds and low latency communication for the delivery of high quality services and analytics.
WiFi Hotspot The smart mirrors can act as wireless access points which would allow other devices to connect to 5G networks through them.
LPWANs Asset Tracking Tracking of the location of physical assets and other personal items would allow users to locate their items quickly and easily.
Energy Management Integration with LPWAN-enabled energy management systems would enable smart mirrors to monitor and adjust energy consumption in real time.
Smart Home Integration Integration with smart home systems that use LPWANs would enable smart mirrors to communicate with other devices in the home, such as smart locks, smart thermostats, and security systems.
Remote Monitoring Users could monitor their mirror’s usage and performance remotely to allow for ease in maintenance and troubleshooting.
The connectivity of the IoM ecosystem is fundamental to its operation and development. To support the envisioned multi-directional communication between smart mirrors and from smart mirror to the user (and vice versa), appropriate network protocols and standards must be considered. Depending on the application and domain, different local and remote wireless communication protocols can be deployed. The connectivity of the IoM would facilitate the processing of computationally heavy commands and requests to off-load the computation from the embedded micro-controller in the smart mirror. Although the applications of the IoM in the cosmetics industry are not time critical in comparison to health applications, real-time use cases such as remote consultancy would still require the network to fulfil low latency, high reliability and synchronisation requirements. The connectivity element of the IoM will not only enable interconnection between IoM smart mirrors and devices but could also facilitate additional services to support the user experience. Table 1 provides an overview of two key wireless technologies, 5G and Low Power Wide Area Networks (LPWANs), and possible features that could be enabled by them in the IoM.
2.3 Applications and Services
Whether the application domain for the IoM is assistive health, beauty or living, there are several services that can be developed from the sensed and communicated data from the smart mirrors and their connectivity features. These services can target various stakeholders including consumers (beauty and cosmetic customers, patients), professionals (cosmeticians, dermatologists, doctors) and service providers (brands, stores, clinics). To establish an efficient interactive interface of the smart mirrors, these services will need to consider real-time computation and requirements. Fig.4 shows the possible remote and co-located interactions between different users in the IoM ecosystem.
@@ 2.3 Applications and Services
2.4 Novel Features
The IoM is envisioned as a unified platform that builds upon and complements existing technologies, designed to work seamlessly with current smart systems while offering enhanced capabilities through interconnectivity. As a framework, the IoM concept presents several novel features that establish its unique value proposition:
  1. 1.
    Unified Health and Beauty Platform: Unlike existing solutions that typically focus on single-purpose or single-domain services, IoM provides a comprehensive platform that recognises and leverages the interconnection between these domains. This integrated approach enables more holistic and personalised recommendations that consider both health and aesthetic factors.
  2. 2.
    Specialised Interconnected Ecosystem: Unlike general IoT systems, the IoM creates a specialised network of smart mirrors designed specifically for health and beauty applications. This mirror-centric approach allows for unique capabilities such as real-time visual analysis, augmented reality overlays for beauty treatments, and a consistent user interface across different locations (home, clinics, retail spaces). The interconnected nature of these mirrors enables a seamless experience as users move between different environments, with their personalised data and preferences following them.
  3. 3.
    Continuous, Non-invasive Monitoring: By leveraging advanced computer vision and AI technologies, the IoM framework complements existing wearable sensors by offering additional non-invasive monitoring capabilities. This approach can augment data from wearables, potentially increasing overall user compliance and providing a more comprehensive dataset for analysis. The combination of wearable data and IoM-derived information offers a richer, more nuanced picture of user health and beauty metrics.
  4. 4.
    Context-aware Personalisation: The IoM system integrates visual data with environmental factors and historical user information to provide highly personalised, context-aware recommendations. This multi-faceted approach to data analysis enables more accurate and relevant suggestions than current standalone systems.
  5. 5.
    Seamless Professional Integration: IoM facilitates easier integration of professional expertise into daily routines. It can serve as a platform for remote consultations with healthcare providers or beauty professionals, bridging the gap between at-home care and professional services.
  6. 6.
    Adaptive Learning: Through its interconnected nature, the IoM system can learn from aggregated, anonymised data across its network. This enables the system to continuously improve its recommendations and adapt to changing trends and user needs.
These novel features address the limitations of current fragmented solutions in health and beauty technologies, as discussed in Section 1.2. By providing a unified, interconnected platform, IoM aims to offer a more comprehensive, personalised, and effective approach to health and beauty care. While these features present significant potential for advancing health and beauty care, it is important to acknowledge that realising the full potential of the IoM framework involves addressing several significant research challenges. The seamless interactions and data exchange across the IoM ecosystem necessitate innovations in various technological domains. These include advanced communication and network technologies, robust security and privacy measures, enhanced interoperability protocols, standardisation efforts, and scalable AI and machine learning algorithms. Addressing these challenges forms a core motivation for the IoM vision, driving forward research and innovation across multiple disciplines. The pursuit of solutions to these open questions may contribute not only to the specific domain of health and beauty care but also to the broader fields of IoT, AI, and connected technologies. Section 5 of this paper provides a more detailed examination of these challenges and potential approaches to address them, underscoring the interdisciplinary nature of the research required to fully realise the IoM framework.
3 Enabling Technologies
In this section, a list of state-of-the-art technologies related to the IoM is presented. We aim to describe the results of various application domains that lead to the conceptualisation of the IoM. A top-level overview of the presented technologies and their potential integration in the IoM is shown at the end of this section in Table 3.
3.1 Standalone Smart Mirrors
@@
In recent years, there have been some increased research efforts within this domain to develop and integrate new features such as the use of Infrared (IR) sensors for user presence detection [37], gesture control and tracking for music control [38] as well as WiFi connectivity [39] for mobile synchronisation with the digitally displayed data [40]. There have also been efforts in the development of plugins and web APIs to support smart mirror applications [41] [42]. There has also been an increase in the number of commercially available smart mirrors for both assistive beauty and assistive living applications. There are many commercial options of stand-alone mirrors for assistive beauty purposes as well as assistive health. Commonly available smart mirrors with their main features have been shown in Table 2.
Table 2: Commercially available smart mirrors and their features.
@@
Sephora Mirror [43] QR code generation for in-store and online makeup, skincare and fragrance recommendations
HiMirror [44] Adjustable lighting, skin analysis, product recommendation, cosmetics tutorial
MemoMi [45] Image capturing and storage,AR eyewear and clothes, Smartphone and SNS sharing
Charlotte Tilbury MagicMirror [46] Augmented makeup, importing images
Shiseido Digital Counseling Mirror [47] Product information, application instructions, skin check, smartphone compatibility
Panasonic New Magic Mirror [48] Skin condition identification, virtual try-on
Modiface [49] Product recommendation, virtual try-on
Lumini Kiosk V2 [50] Skin analysis, product recommendation, skin comparison with other users
Icon AI Z Mirror [51] Adjustable lighting, skin analysis, product recommendation, virtual try-on
Morror [52] Adjustable lighting, cosmetics tutorial, weather information
Simplehuman Sensor Mirror Pro [53] Smart lighting
Kohler Verdera Voice Lighted Mirror [54] Voice activation integrated with Alexa and Google assistant
iHome Reflect Smart Vanity Mirror [55] Lighting and speaker control with smartphone
Capstone Connected Home Smart [56] Integration with other smart home devices, built-in speaker, voice activation
@@
Figure 5: A simple IoT-enabled smart mirror schematic.
3.2 Skin Imaging Methodologies
@@
3.3 Real-Time Health Monitoring